How much must we invest at the end of each month to build a balance of $610,000 over 14 years if we earn 9% per annum? (Round to nearest penny and enter, for example, as 123456.78) (Can you please also show how to work it out on financial calculator BA II plus if possible)
Answer #1

Calculate for PMT, punch in:-

I/Y =9%/12 =0.75%

N =14*12 =168

FV =610000

The answer will be 1823.52